OpenStudy (anonymous):

Factor completely: 21x3 + 35x2 + 9x + 15

OpenStudy (anonymous):

(3x − 5)(7x^2 − 3) = (3x)(7x^2) + (3x)(-3) + (-5)(7x^2) + (-5)(-3) = 21x^3 - 9x - 35x^2 +15 = 21x^3 - 35x^2 - 9x + 15. Not equal. (3x − 5)(7x2 + 3) = 21x^3 + 9x - 35x^2 ... Not equal. (3x + 5)(7x2 − 3) = 21x^3 - 9x ... Not equal.

OpenStudy (anonymous):

=7x2(3x+5)+3(3x+5)=(3x+5)(7x2+3)

OpenStudy (apoorvk):

basically you need hit and trial to find any one factor of the poly, then you can use synthetic division by (x-factor) to get a quadratic, factorising which you will get all three roots and hence the factors.

OpenStudy (apoorvk):

or if you can directly factorise like a quadratic like arnab has done above, wonderful too!! (though doesnt happen every time)

OpenStudy (anonymous):

factoring in groups: help.... 7x^3+5x^2-21x-15=